# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ D.4 Mirror scheme with non-pushforward compatible Bregman potentials
We study in this Section schemes for which the Bregman potential $\phi$ is not pushforward compatible, and thus for which we cannot apply Proposition 1 and thus 1 does not hold a priori. An example of such potential is $\phi_{\mu}(\mathrm{T})=\langle\mathrm{T},P_{\mu}\mathrm{T}\rangle_{L^{2}(\mu)}$ where $P_{\mu}:L^{2}(\mu)\to L^{2}(\mu)$ is a linear autoadjoint and invertible operator. Since $\nabla\phi_{\mu}(\mathrm{T})=P_{\mu}\mathrm{T}$, taking the first order conditions, we obtain the following scheme:
@@
In particular, this includes SVGD [77, 76, 66] (if we pose $P_{\mu}^{-1}\mathrm{T}=\iota S_{\mu}\mathrm{T}$ with $S_{\mu}:L^{2}(\mu)\to\mathcal{H}$ defined as $S_{\mu}\mathrm{T}=\int k(x,\cdot)\mathrm{T}(x)\mathrm{d}\mu(x)$ which maps functions from $L^{2}(\mu)$ to the reproducing kernel Hilbert space $\mathcal{H}$ with kernel $k$, and with $\iota:\mathcal{H}\to L^{2}(\mu)$ the inclusion operator that is the adjoint of $S_{\mu}$ [66]) or the Kalman-Wasserstein gradient descent [52] (for which $P_{\mu}^{-1}=\int\big{(}x-m(\mu)\big{)}\otimes\big{(}x-m(\mu)\big{)}\ \mathrm{d}\mu(x)$ is the covariance matrix, where $m(\mu)=\int x\ \mathrm{d}\mu(x)$).
@@
$\forall\mathrm{T},\mathrm{S}\in L^{2}(\!u),\ \mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\frac{1}{2}\langle\mathrm{T},C(\mu)\mathrm{T}\rangle_{L^{2}(\mu)}+\frac{1}{2}\langle\mathrm{S},C(\mu)\mathrm{S}\rangle_{L^{2}(\mu)}-\langle C(\mu)\mathrm{S},\mathrm{T}\rangle_{L^{2}(\mu)}$ (77)
@@
Moreover, the problem $\inf_{\gamma\in\Pi(\alpha,\beta)}\ \int\|C(\mu)^{\frac{1}{2}}(x-y)\|_{2}^{2}\ \mathrm{d}\gamma(x,y)$ is equivalent with
@@
where we assume that $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\in L^{2}(\mu)$. In that case, using the first order conditions,
@@ E.1 Relative convexity and smoothness between Fenchel transforms
$\bar{f}\in\partial\phi^{*}(g)\iff\phi^{*}(\bar{f})=\langle\bar{f},g\rangle-\phi(\bar{f})=\sup_{f\in L^{2}(\mu)}\langle f,g\rangle-\phi(f)$
So to prove there is a unique element in $\partial\phi^{*}(g)$, we just need to show that, setting $\phi_{g}(f):=-\langle f,g\rangle+\phi(f)$, the problem $\inf_{f\in L^{2}(\mu)}\phi_{g}(f)$ has a unique solution. Because of our assumptions, $\phi_{g}$ is lower semi continuous and strictly convex. Since $\phi$ is superlinear, $\phi_{g}$ is coercive, i.e. $\lim_{\|f\|\to\infty}\phi_{g}(f)=+\infty$. There thus exists a solution [7, Theorem 3.3.4], which is unique by strict convexity. Hence $\partial\phi^{*}(g)$ is reduced to a point, which is necessarily the Gâteaux derivative of $\phi^{*}$ at $g$. ∎
This allows us to relate the Bregman divergence of $\phi^{*}$ with the Bregman divergence of $\phi$.
@@
By [92, Corollary 3.44], we have $\phi^{*}\big{(}\nabla\phi(\mathrm{T})\big{)}=\langle\mathrm{T},\nabla\phi(\mathrm{T})\rangle_{L^{2}(\mu)}-\phi(\mathrm{T})$ for all $\mathrm{T}\in L^{2}(\mu)$ since $\phi$ is convex and differentiable. By Lemma 16, $\phi^{*}$ is invertible and by [9, Corollary 16.24], since $\phi$ is proper, lower semi-continuous and convex, then $(\nabla\phi)^{-1}=\nabla\phi^{*}$.
@@
$=\langle\nabla\phi(\mathrm{T}),\mathrm{T}\rangle_{L^{2}(\mu)}-\phi(\mathrm{T})-\langle\nabla\phi(\mathrm{S}),\mathrm{S}\rangle_{L^{2}(\mu)}+\phi(\mathrm{S})$
@@
$=\mathrm{d}_{\phi}(\mathrm{S},\mathrm{T}).$ (82)
Finally, we can relate the relative convexity of $\phi$ relative to $\psi^{*}$ by using an inequality between the Bregman divergences of $\phi$ and $\psi$. In particular, we recover the assumptions of Propositions 4 and 5 for $\phi_{\mu_{k}}^{h^{*}}$ $\beta$-smooth and $\alpha$-convex relative to ~$\tilde{\mathcal{F}}_{\mu_{k}}^{*}$.
@@
Let $\phi,\psi:L^{2}(\mu)\to\mathbb{R}$ proper, superlinear, strictly convex and differentiables. $\phi$ is $\beta$-smooth (resp. $\alpha$-convex) relative to $\psi^{*}$ if and only if $\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, $\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}\leq\beta\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T})$ (resp. $\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}\geq\alpha\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T})$).
@@
$\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}$ $\leq\beta\mathrm{d}_{\psi^{*}}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}$ (84)
$=\beta\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T}),$
using Lemma 17 in the last line.
On the other hand, by Lemma 17, for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu),$
$\mathrm{d}_{\psi^{*}}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}=\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T})\geq\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}.$ (85)
@@
Proposition 18 allows to translate the inequality hypothesis of Proposition 4 and Proposition 5. Assume that for all $k$, ~$\tilde{\mathcal{F}}_{\mu_{k}}$ is strictly convex, differentiable and superlinear. We note first that it implies that $\mathcal{F}_{\mu_{k}}$ is convex along $t\mapsto\big{(}(1-t)\mathrm{T}_{k+1}+t\mathrm{Id}\big{)}_{\#}\mu_{k}$. Moreover, by Lemma 16, ~$\nabla\tilde{\mathcal{F}}_{\mu_{k}}^{*}$ is differentiable.
@@
For a curve $t\mapsto\mu_{t}$, we define $\mathcal{F}_{\mu}^{*}$ on $\mu_{t}$ as $\mathcal{F}_{\mu}^{*}(\mu_{t}):=\tilde{\mathcal{F}}_{\mu}^{*}(\mathrm{T}_{t})$ with ~$\tilde{\mathcal{F}}_{\mu}^{*}$ the convex conjugate of ~$\tilde{\mathcal{F}}_{\mu}$ in the $L^{2}(\mu)$ sense. Then, we can apply Proposition 18, and we obtain that the inequality hypothesis of Proposition 4 is equivalent to the $\beta$-smoothness of $\phi^{h^{*}}$ relative to $\mathcal{F}_{\mu_{k}}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\circ\mathrm{T}_{k+1}\big{)}_{\#}\mu_{k}$ since
@@
$=\beta\mathrm{d}_{\tilde{\mathcal{F}}^{*}_{\mu_{k}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\circ\mathrm{T}_{k+1},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}.$
@@
is equivalent to the $\alpha$-convexity of $\phi^{h^{*}}$ relative to $\mathcal{F}_{\mu_{k}}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu_{k})\circ\mathrm{T}\big{)}_{\#}\mu_{k}$.
@@
Let $(X,\tau)$ be a metrizable topological space, and $f:X\to\mathbb{R}\cup\{+\infty\}$ be strictly convex, $\tau$-lower semicontinuous and with one $\tau$-compact sublevel set. Let $x_{0}\in X$ be the minimizer of $f$ and take a sequence $(x_{n})_{n\in\mathbb{N}}$ such that $f(x_{n})\to f(x_{0})$ then $(x_{n})_{n\in\mathbb{N}}$ $\tau$-converges to $x_{0}$.
@@
The existence of the minimum is given by [7, Theorem 3.2.2]. For $N$ large enough, $(x_{n})_{n\geq N}$ lives in the $\tau$-compact sublevel set, since $x_{0}$ belongs to it and $f(x_{0})$ is minimal. We can then consider a subsequence $\tau$-converging to some $x^{*}$. By $\tau$-lower semicontinuity, we have $f(x_{0})\leq f(x^{*})\leq\liminf f(x_{\sigma(n)})=f(x_{0})$, so $f(x_{0})=f(x^{*})$ and by strict convexity $x_{0}=x^{*}$. Since all subsequences of $(x_{n})_{n\geq N}$ converge to $x^{*}$ and the space is metrizable, $(x_{n})_{n\in\mathbb{N}}$ $\tau$-converges to $x_{0}$. ∎
The typical case is when $X$ is a Hilbert space and $\tau$ is the weak topology. One could wish to have strong convergence under a coercivity assumption, however “In infinite dimensional spaces, the topologies which are directly related to coercivity are the weak topologies” [7, p86]. Nevertheless Gâteaux differentiability implies continuity, which paired with convexity gives weak lower semicontinuity [7, Theorem 3.3.3]. We cannot hope for convergence of the norm of $x_{n}$ to come for free, as the weak convergence would then imply the strong convergence.
@@ E.2 Relative convexity and smoothness between functionals
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\phi_{\mu}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ and $\psi_{\mu}(\mathrm{T})=\int U\circ\mathrm{T}\ \mathrm{d}\mu$ where $V:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{1}$. If $V$ is $\alpha$-convex (resp. $\beta$-smooth) relative to $U:\mathbb{R}^{d}\to\mathbb{R}$, then $\phi_{\mu}$ is $\alpha$-convex (resp $\beta$-smooth) relative to $\psi_{\mu}$.
@@
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $W,K:\mathbb{R}^{d}\to\mathbb{R}$ be symmetric, $C^{1}$ and convex. Let $\phi_{\mu}(\mathrm{T})=\iint W\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ and $\psi_{\mu}(\mathrm{T})=\iint K\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$. If $W$ is $\alpha$-convex relative to $K$, then $\phi_{\mu}$ is $\alpha$-convex relatively to $\psi_{\mu}$. Likewise, if $W$ is $\beta$-smooth relatively to $K$, then $\phi_{\mu}$ is $\beta$-smooth relatively to $\psi_{\mu}$.
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\iint\mathrm{d}_{W}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ (98)
$\geq\alpha\iint\mathrm{d}_{K}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$
@@
Thus, in situations where the objective functional and the Bregman potential are of the same type and either potential energies or interaction energies, we only need to show the convexity and smoothness of the underlying potentials or interaction kernels. For instance, let $V:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function, such that $\|\nabla^{2}V\|_{\mathrm{op}}\leq p_{r}(\|x\|_{2})$ with $p_{r}$ a polynomial function of degree $r$ and $\|\cdot\|_{\mathrm{op}}$ the operator norm. Then, by [81, Proposition 2.1], $V$ is $\beta$-smooth relative to $h$ where for all $x\in\mathbb{R}^{d}$, $h(x)=\frac{1}{r+2}\|x\|_{2}^{r+2}+\frac{1}{2}\|x\|_{2}^{2}$.
@@
along the right $\mu_{t}$.
For instance, suppose $\mathcal{F}(\mu)=\frac{1}{2}\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ and $\mathcal{G}(\mu)=\int V\mathrm{d}\mu$. Then, by Example 1 and Example 2, we have, for $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$ and $\mathrm{T}_{t}=\mathrm{S}+tv$,
@@
To show the conditions of Proposition 2, we need to take $\mathrm{S}=\mathrm{Id}$ and $v=\mathrm{T}_{k+1}-\mathrm{Id}$, and to verify for $t=0$ the inequality, i.e.
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{F}(\mu_{t})\Big{|}_{t=0}\leq\beta\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{G}(\mu_{t})\Big{|}_{t=0}$ (102)
@@
$\iint\langle\nabla^{2}W(x-y)\big{(}v(x)-v(y)\big{)},v(x)\rangle\ \mathrm{d}\mu_{k}(x)\mathrm{d}\mu_{k}(y)\leq\beta\int\langle\nabla^{2}V(x)v(x),v(x)\rangle\ \mathrm{d}\mu_{k}(x)$

### v2

@@ D.4 Mirror scheme with non-pushforward compatible Bregman potentials
We study in this Section schemes for which the Bregman potential $\phi_{\mu}$ is not pushforward compatible, and thus for which we cannot apply Proposition 2 and thus 1 may not hold a priori. An example of such potential is $\phi_{\mu}(\mathrm{T})=\langle\mathrm{T},P_{\mu}\mathrm{T}\rangle_{L^{2}(\mu)}$ where $P_{\mu}:L^{2}(\mu)\to L^{2}(\mu)$ is a linear autoadjoint and invertible operator. Since $\nabla\phi_{\mu}(\mathrm{T})=P_{\mu}\mathrm{T}$, taking the first order conditions, we obtain the following scheme:
@@
In particular, this includes SVGD [84, 83, 71] if we pose $P_{\mu}^{-1}\mathrm{T}=\iota S_{\mu}\mathrm{T}$ with $S_{\mu}:L^{2}(\mu)\to\mathcal{H}$ defined as $S_{\mu}\mathrm{T}=\int k(x,\cdot)\mathrm{T}(x)\mathrm{d}\mu(x)$ which maps functions from $L^{2}(\mu)$ to the reproducing kernel Hilbert space $\mathcal{H}$ with kernel $k$, and with $\iota:\mathcal{H}\to L^{2}(\mu)$ the inclusion operator that is the adjoint of $S_{\mu}$ [71]. It also includes the Kalman-Wasserstein gradient descent [56] for which $P_{\mu}^{-1}=\int\big{(}x-m(\mu)\big{)}\otimes\big{(}x-m(\mu)\big{)}\ \mathrm{d}\mu(x)$ is the covariance matrix, where $m(\mu)=\int x\ \mathrm{d}\mu(x)$.
@@
$\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\frac{1}{2}\langle\mathrm{T},C(\mu)\mathrm{T}\rangle_{L^{2}(\mu)}+\frac{1}{2}\langle\mathrm{S},C(\mu)\mathrm{S}\rangle_{L^{2}(\mu)}-\langle C(\mu)\mathrm{S},\mathrm{T}\rangle_{L^{2}(\mu)}$ (77)
@@
Moreover, the problem $\inf_{\gamma\in\Pi(\alpha,\beta)}\ \int\|C(\mu)^{\frac{1}{2}}(x-y)\|_{2}^{2}\ \mathrm{d}\gamma(x,y)$ is equivalent to
@@
where we assume that $\nabla_{\mathrm{W}}\mathcal{F}(\mu)\in L^{2}(\mu)$. In that case, using the first order conditions,
@@ E.1 Relative convexity and smoothness between Fenchel transforms
$\bar{f}\in\partial\phi^{*}(g)\iff\phi^{*}(g)=\langle\bar{f},g\rangle-\phi(\bar{f})=\sup_{f\in L^{2}(\mu)}\langle f,g\rangle-\phi(f).$ (82)
So to prove there is a unique element in $\partial\phi^{*}(g)$, we just need to show that, setting $\phi_{g}(f):=-\langle f,g\rangle+\phi(f)$, the problem $\inf_{f\in L^{2}(\mu)}\phi_{g}(f)$ has a unique solution. Under our assumptions, $\phi_{g}$ is lower semicontinuous and strictly convex. Since $\phi$ is superlinear, $\phi_{g}$ is coercive, i.e. $\lim_{\|f\|\to\infty}\phi_{g}(f)=+\infty$. There thus exists a solution [8, Theorem 3.3.4], which is unique by strict convexity. Hence $\partial\phi^{*}(g)$ is reduced to a point, which is necessarily the Gâteaux derivative of $\phi^{*}$ at $g$. ∎
This allows us to relate the Bregman divergence of $\phi^{*}$ to the Bregman divergence of $\phi$.
@@
By [99, Corollary 3.44], we have $\phi^{*}\big{(}\nabla\phi(\mathrm{T})\big{)}=\langle\mathrm{T},\nabla\phi(\mathrm{T})\rangle_{L^{2}(\mu)}-\phi(\mathrm{T})$ for all $\mathrm{T}\in L^{2}(\mu)$ since $\phi$ is convex and differentiable. By Lemma 18, $\phi^{*}$ is invertible and by [10, Corollary 16.24], since $\phi$ is proper, lower semicontinuous and convex, then $(\nabla\phi)^{-1}=\nabla\phi^{*}$.
@@
$=\langle\nabla\phi(\mathrm{T}),\mathrm{T}\rangle_{L^{2}(\mu)}-\phi(\mathrm{T})-\langle\nabla\phi(\mathrm{S}),\mathrm{S}\rangle_{L^{2}(\mu)}+\phi(\mathrm{S})$ (83)
@@
$=\mathrm{d}_{\phi}(\mathrm{S},\mathrm{T}).\qed$
Finally, we can relate the relative convexity of $\phi$ relative to $\psi^{*}$ by using an inequality between the Bregman divergences of $\phi$ and $\psi$. In particular, we recover the assumptions of Propositions 5 and 6 for $\phi_{\mu_{k}}^{h^{*}}$ that is $\beta$-smooth and $\alpha$-convex relative to ~$\tilde{\mathcal{F}}_{\mu_{k}}^{*}$.
@@
Let $\phi,\psi:L^{2}(\mu)\to\mathbb{R}$ proper, superlinear, strictly convex and differentiable. $\phi$ is $\beta$-smooth (resp. $\alpha$-convex) relative to $\psi^{*}$ if and only if $\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, $\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}\leq\beta\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T})$ (resp. $\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}\geq\alpha\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T})$).
@@
$\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}$ $\leq\beta\mathrm{d}_{\psi^{*}}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}=\beta\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T}),$ (85)
using Lemma 19.
On the other hand, suppose for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, $\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}\leq\beta\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T})$. Then, by first using Lemma 19 and then the supposed inequality, we have for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu),$
$\beta\mathrm{d}_{\psi^{*}}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}=\beta\mathrm{d}_{\psi}(\mathrm{S},\mathrm{T})\geq\mathrm{d}_{\phi}\big{(}\nabla\psi(\mathrm{T}),\nabla\psi(\mathrm{S})\big{)}.$ (86)
@@
Proposition 20 allows to translate the inequality hypothesis of Proposition 5 and Proposition 6. Assume that for all $k$, ~$\tilde{\mathcal{F}}_{\mu_{k}}$ is strictly convex, differentiable and superlinear. We first note that it implies that $\mathcal{F}_{\mu_{k}}$ is convex along $t\mapsto\big{(}(1-t)\mathrm{T}_{k+1}+t\mathrm{Id}\big{)}_{\#}\mu_{k}$. Moreover, by Lemma 18, ~$\nabla\tilde{\mathcal{F}}_{\mu_{k}}^{*}$ is differentiable.
@@
For a curve $t\mapsto\mu_{t}$, we define $\mathcal{F}_{\mu}^{*}$ on $\mu_{t}$ as $\mathcal{F}_{\mu}^{*}(\mu_{t}):=\tilde{\mathcal{F}}_{\mu}^{*}(\mathrm{T}_{t})$ with ~$\tilde{\mathcal{F}}_{\mu}^{*}$ the convex conjugate of ~$\tilde{\mathcal{F}}_{\mu}$ in the $L^{2}(\mu)$ sense. Then, we can apply Proposition 20, and we obtain that the inequality hypothesis of Proposition 5 is implied by the $\beta$-smoothness of $\phi^{h^{*}}$ relative to $\mathcal{F}_{\mu_{k}}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\circ\mathrm{T}_{k+1}\big{)}_{\#}\mu_{k}$ since
@@
$=\beta\mathrm{d}_{\tilde{\mathcal{F}}^{*}_{\mu_{k}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\circ\mathrm{T}_{k+1},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)},$
where we used Proposition 1 to compute the gradient ~$\nabla\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{T}_{k+1})=\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\circ\mathrm{T}_{k+1}$.
@@
is implied by the $\alpha$-convexity of $\phi^{h^{*}}$ relative to $\mathcal{F}_{\mu_{k}}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu_{k})\circ\mathrm{T}\big{)}_{\#}\mu_{k}$.
These results are summarized in Proposition 7 and shown formally in Section H.7.
@@
Let $(X,\tau)$ be a metrizable topological space, and $f:X\to\mathbb{R}\cup\{+\infty\}$ be strictly convex, $\tau$-lower semicontinuous and with one $\tau$-compact sublevel set. Let $x_{0}\in X$ be the minimizer of $f$ and take a sequence $(x_{n})_{n\in\mathbb{N}}$ such that $f(x_{n})\to f(x_{0})$. Then, $(x_{n})_{n\in\mathbb{N}}$ $\tau$-converges to $x_{0}$.
@@
The existence of the minimum is given by [8, Theorem 3.2.2]. For $N$ large enough, $(x_{n})_{n\geq N}$ lives in the $\tau$-compact sublevel set of $f$, since $x_{0}$ belongs to it and $f(x_{0})$ is minimal. We can then consider a subsequence $\tau$-converging to some $x^{*}$. By $\tau$-lower semicontinuity, we have $f(x_{0})\leq f(x^{*})\leq\liminf f(x_{\sigma(n)})=f(x_{0})$, so $f(x_{0})=f(x^{*})$ and by strict convexity $x_{0}=x^{*}$. Since all subsequences of $(x_{n})_{n\geq N}$ converge to $x^{*}$ and the space is metrizable, $(x_{n})_{n\in\mathbb{N}}$ $\tau$-converges to $x_{0}$. ∎
The typical case is when $X$ is a Hilbert space and $\tau$ is the weak topology. One could wish to have strong convergence under a coercivity assumption, however “In infinite dimensional spaces, the topologies which are directly related to coercivity are the weak topologies” [8, p86]. Nevertheless Gâteaux differentiability implies continuity, which paired with convexity gives weak lower semicontinuity [8, Theorem 3.3.3]. We cannot hope for convergence of the norm of $x_{n}$ to come for free, as the weak convergence would then imply the strong convergence.
@@ E.2 Relative convexity and smoothness between functionals
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\phi_{\mu}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ and $\psi_{\mu}(\mathrm{T})=\int U\circ\mathrm{T}\ \mathrm{d}\mu$ where $V,U:\mathbb{R}^{d}\to\mathbb{R}$ are $C^{1}$. If $V$ is $\alpha$-convex (resp. $\beta$-smooth) relative to $U:\mathbb{R}^{d}\to\mathbb{R}$, then $\phi_{\mu}$ is $\alpha$-convex (resp $\beta$-smooth) relative to $\psi_{\mu}$.
@@
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $W,K:\mathbb{R}^{d}\to\mathbb{R}$ be symmetric, $C^{1}$ and convex. Let $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\iint W\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ and $\psi_{\mu}(\mathrm{T})=\frac{1}{2}\iint K\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$. If $W$ is $\alpha$-convex relative to $K$, then $\phi_{\mu}$ is $\alpha$-convex relatively to $\psi_{\mu}$. Likewise, if $W$ is $\beta$-smooth relatively to $K$, then $\phi_{\mu}$ is $\beta$-smooth relatively to $\psi_{\mu}$.
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\frac{1}{2}\iint\mathrm{d}_{W}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ (99)
$\geq\frac{\alpha}{2}\iint\mathrm{d}_{K}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$
@@
Thus, in situations where the objective functional and the Bregman potential are of the same type and either potential energies or interaction energies, we only need to show the convexity and smoothness of the underlying potentials or interaction kernels. For instance, let $V:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function, such that $\|\nabla^{2}V\|_{\mathrm{op}}\leq p_{r}(\|x\|_{2})$ with $p_{r}$ a polynomial function of degree $r$ and $\|\cdot\|_{\mathrm{op}}$ the operator norm. Then, by [88, Proposition 2.1], $V$ is $\beta$-smooth relative to $h$ where for all $x\in\mathbb{R}^{d}$, $h(x)=\frac{1}{r+2}\|x\|_{2}^{r+2}+\frac{1}{2}\|x\|_{2}^{2}$.
@@
along the right curve $t\mapsto\mu_{t}$.
For instance, consider an objective functional $\mathcal{F}(\mu)=\frac{1}{2}\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ and another functional $\mathcal{G}(\mu)=\int V\mathrm{d}\mu$. Then, by Example 1 and Example 2, we have, for $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$ and $\mathrm{T}_{t}=\mathrm{S}+tv$,
@@
To show the conditions of Proposition 3, we need to take $\mathrm{S}=\mathrm{Id}$ and $v=\mathrm{T}_{k+1}-\mathrm{Id}$, and to verify for $t=s\in[0,1]$ the inequality, i.e.
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{F}(\mu_{t})\Big{|}_{t=s}\leq\beta\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{G}(\mu_{t})\Big{|}_{t=s}$ (103)
@@
$\iint\langle\nabla^{2}W\big{(}\mathrm{T}_{s}(x)-\mathrm{T}_{s}(y)\big{)}\big{(}v(x)-v(y)\big{)},v(x)\rangle\ \mathrm{d}\mu_{k}(x)\mathrm{d}\mu_{k}(y)$
$\leq\beta\int\langle\nabla^{2}V\big{(}\mathrm{T}_{s}(x)\big{)}v(x),v(x)\rangle\ \mathrm{d}\mu_{k}(x)$